OpenStudy (anonymous):

how can you say on the word's of physics. what is real definition of electric current and how are chage is created ?

OpenStudy (anonymous):

Physics is best expressed through mathematics. Current is \[I(t) = dq/dt = d/dt \int\limits_{}dq dA\] It's how much charge passes through a given area, in a given amount of time. It's a change in current, per change in time. You can change current by changing the force that is applied to the flowing electrons via some force/potential.

OpenStudy (anonymous):

er... I said "change in current per change in time" but I meant to say, change in charge per change in time.